OpenStudy (kristy1234567):

What are the x and y intercepts of the equation? 4x−3y=12 A: x = 3 and y = 4 B: x=−4 and y = 3 C: x=−4 and y=−3 D: x = 3 and y=−4

OpenStudy (amit1235813):

For x intercept, plug y = 0

OpenStudy (helpppp):

Fill in for x and y

OpenStudy (kristy1234567):

i hate math its hard DX i dont learn the same way as other ppl

OpenStudy (amit1235813):

If y = 0, what do you get?

OpenStudy (kristy1234567):

i dont know DX

OpenStudy (helpppp):

4x-3(0)=12 4x-0=12 12/4=x

OpenStudy (kristy1234567):

THAT makes no sense to me

OpenStudy (kristy1234567):

sorry :( i just dont understand math its my worst subject i failed math last year

OpenStudy (helpppp):

Do you have an email. My openstudy is being crap rn

OpenStudy (tyrenhere):

You could just plug in. Try every answer.

OpenStudy (kristy1234567):

yes i do

OpenStudy (tyrenhere):

Even though I already told you it's A

OpenStudy (kristy1234567):

it is? before u told me something with a 6 or something sorry i dont pay attention much in math DX its my worst subject

OpenStudy (helpppp):

@TyrenHere Thanks for just giving away the answer. :P We are supposed to help them understand how to do the problem, if not they will just keep asking

OpenStudy (tyrenhere):

I had explained it to her already in math chat Lol

OpenStudy (kristy1234567):

i sent u my email helpppp

OpenStudy (tyrenhere):

Odam

OpenStudy (tyrenhere):

Uhm. I don't have my calculator. Just try all for solutions